OpenStudy (anonymous):

solve sin^2(theta)=2sin(theta) +3

OpenStudy (anonymous):

If you take x = sin(theta) you could rewrite that as x^2 = 2x + 3 and then solve for x and plug back for sin(theta)

OpenStudy (anonymous):

so x will equal 3,-1?

OpenStudy (anonymous):

Yup, now you just have to plug back in sin(theta) and get theta all by it's self.

OpenStudy (anonymous):

so itll be 3^2=2(3)+3 like that?

OpenStudy (anonymous):

Well you already have x = 3 and since you substituted sin(theta) for x you still need to find out what theta is You have right now that x = 3 and -1 so then sin(theta) would also = 3 and -1

OpenStudy (anonymous):

so the answer will be theta = 3, -1

OpenStudy (anonymous):

well you still have sin(theta), you can not just get rid of the sine function. If you use algebra arcsin(sin(theta)) = theta So you just have to use that logic to solve for theta in the equations you already have. sin(theta) = 3 & sin(theta) = -1

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Your welcome!